# Quenched large deviations of Birkhoff sums along random quantum measurements

### v1

@@ 1 Introduction
$p^{(n)}(a_{1},\dotsc,a_{n})=\operatorname{tr}[(\psi_{n,a_{n}}\circ\dotsb\circ\psi_{1,a_{1}})(\rho)].$
@@ 2 Setup and main result
$M_{\omega,\rho}^{(n)}(\alpha)=\operatorname{tr}\left[\left(\varphi_{\theta^{n-1}\omega}^{(\alpha)}\circ\cdots\circ\varphi_{\omega}^{(\alpha)}\right)(\rho)\right].$ (2.3)
@@
For every fixed $\alpha$, the results of [MS22, PS23] will apply and tell us — among other things — that there is an $\mathbb{S}^{\circ}$-valued random variable $Z^{(\alpha)}$ that satisfies the cocycle relation
$\varphi_{\omega}^{(\alpha)}\left(Z_{\theta^{-1}\omega}^{\alpha}\right)=Z_{\omega}^{(\alpha)}\operatorname{tr}\left[\left(Z_{\theta^{-1}\omega}^{\alpha}\right)\right]$
@@
These objects are at the heart of the proof of Theorem 2.3, which has of three main steps and is carried out in Section 4. We summarize the main steps of the proof here:
@@
In view of the specific form of the dependence of $\varphi_{\omega}^{(\alpha)}$ in $\alpha$, we use arguments from [MS22, PS23] to show that there is an event of full measure on which the limiting random variables $Z^{(\alpha)}$ (and analogous objects for adjoints) and the top Lyapunov exponent $\lambda(\alpha)$ exist simultaneously for every deformation parameter $\alpha\in{\mathbb{R}}$. This is the content of Lemmas 4.6 and 4.8, below.
@@ 3 Application to entropy production
Suppose that we have measurable maps $\omega\mapsto U_{\omega}\in\operatorname{U}({\mathbb{C}}^{m})$ and $\omega\mapsto\xi_{\omega}$, where $\xi_{\omega}$ a Gibbs state on ${\mathbb{C}}^{m}$ with spectral decomposition
@@
using rank-one projectors, increasing energy labeling $\epsilon\mapsto E_{\omega,\epsilon}$ and positive inverse temperature $\beta_{\omega}$. Then, consider instruments labeled by pairs $(\epsilon,\epsilon^{\prime})\in\{1,2,\dotsc,m\}^{2}$ of indices for the spectral decomposition, that are of the form
@@
$f_{\omega}(\epsilon,\epsilon^{\prime})=\beta_{\omega}(E_{\omega,\epsilon^{\prime}}-E_{\omega,\epsilon}).$
@@
$\Sigma_{n}((\epsilon_{1},\epsilon^{\prime}_{1}),(\epsilon_{2},\epsilon^{\prime}_{2}),\dotsc,(\epsilon_{n},\epsilon^{\prime}_{n}))=\sum_{j=0}^{n-1}\beta_{\theta^{j}\omega}(E_{{\theta^{j}\omega},\epsilon^{\prime}_{j}}-E_{{\theta^{j}\omega},\epsilon_{j}}).$
@@
and antiunitaries $\tau:{\mathbb{C}}^{d}\to{\mathbb{C}}^{d}$ and $\tau^{\prime}:{\mathbb{C}}^{m}\to{\mathbb{C}}^{m}$ such that
@@
    In the setup of Example 2.1.i, the map $\mathsf{T}:(\omega_{n})_{n\in{\mathbb{Z}}}\mapsto(\omega_{-n})_{n\in{\mathbb{Z}}}$ clearly satisfies the condition on the interplay of $\mathsf{T}$ and $\theta$, and the condition on the interplay of $\mathbb{P}$ and $\mathsf{T}$ is then naturally interpreted as a generalization of detailed balance for Markov chains. With $\omega\mapsto U_{\omega}\in\operatorname{U}({\mathbb{C}}^{m})$ and $\omega\mapsto\xi_{\omega}$ depending only on $\omega_{0}$, the conditions on $\tau$ and $\tau^{\prime}$ further simplify to $(\tau\otimes\tau^{\prime})U_{\omega}=U^{*}_{\omega}(\tau\otimes\tau^{\prime})$ and $\tau^{\prime}\pi_{\omega,\epsilon}=\pi_{\omega,\epsilon}\tau^{\prime}$. Writing $U_{\omega}=\exp(\mathrm{i}t_{\omega}H_{\omega})$, note that such $\tau$ and $\tau^{\prime}$ exist as complex conjugations if there exist orthonormal bases of ${\mathbb{C}}^{d}$ and ${\mathbb{C}}^{m}$ (used to construct an orthonormal basis of ${\mathbb{C}}^{d}\otimes{\mathbb{C}}^{m}$) with respect to which $\xi_{\omega}$ and $H_{\omega}$ have real matrix elements.
@@
Fix $\omega$ and suppose that $\xi_{\omega}=\tau^{\prime}\xi_{\mathsf{T}\omega}\tau^{\prime}$ as required in TRI. While we have already chosen a spectral decomposition $\xi_{\omega}=\sum_{\epsilon}\mathrm{e}^{-\beta_{\omega}E_{\omega,\epsilon}}\pi_{\omega,\epsilon}$ for the left-hand side of the equality (and for its analogue $\xi_{\mathsf{T}\omega}=\sum_{\epsilon}\mathrm{e}^{-\beta_{\mathsf{T}\omega}E_{\mathsf{T}\omega,\epsilon}}\pi_{\mathsf{T}\omega,\epsilon}$), we have some freedom in choosing the spectral decomposition of the right-hand side of the equality. Suppose that
$\tau^{\prime}\xi_{\mathsf{T}\omega}\tau^{\prime}=\sum_{{\epsilon}=1}^{n}\mathrm{e}^{-\tilde{Q}_{{\epsilon}}}\tilde{\pi}_{{\epsilon}}$
is also indexed so that $\tilde{Q}_{1}<\tilde{Q}_{2}<\dotsb<\tilde{Q}_{m}$. Since the eigenvalues of $\tau^{\prime}\xi_{\mathsf{T}\omega}\tau^{\prime}$ and $\xi_{\mathsf{T}\omega}$ coincide, it must be the case that $\tilde{Q}_{{\epsilon}}=\beta_{\mathsf{T}\omega}E_{\mathsf{T}{\omega},\epsilon}$. But then, since $\tau^{\prime}\xi_{\mathsf{T}\omega}\tau^{\prime}$ can also be written as $\sum_{\epsilon}\mathrm{e}^{-\beta_{\mathsf{T}\omega}E_{\mathsf{T}\omega,\epsilon}}\tau^{\prime}\pi_{\mathsf{T}\omega,\epsilon}\tau^{\prime}$, it must be the case that $\tau^{\prime}\pi_{\mathsf{\omega},\epsilon}\tau^{\prime}=\tilde{\pi}_{\epsilon}$. Going back to the equality we started with, and using yet again the fact that the eigenvalues are ordered, this means that $\tau^{\prime}\pi_{\mathsf{\omega},\epsilon}\tau^{\prime}=\pi_{\mathsf{T}\omega,\epsilon}$.
@@
$\sigma_{\omega,\rho,n}((\epsilon_{1},\epsilon^{\prime}_{1}),(\epsilon_{2},\epsilon^{\prime}_{2}),\dotsc,(\epsilon_{n},\epsilon^{\prime}_{n})):=\log\frac{p^{(n)}_{\omega,\rho}((\epsilon_{1},\epsilon^{\prime}_{1}),(\epsilon_{2},\epsilon^{\prime}_{2}),\dotsc,(\epsilon_{n},\epsilon^{\prime}_{n}))}{p^{(n)}_{\mathsf{T}(\theta^{n-1}\omega),\rho}((\epsilon^{\prime}_{n},\epsilon_{n}),(\epsilon^{\prime}_{n-1},\epsilon_{n-1}),\dotsc,(\epsilon^{\prime}_{1},\epsilon_{1}))}.$ (3.1)
@@
$\Delta_{\rho}=\frac{1}{\inf\operatorname{sp}\rho}.$
@@ 4.1 Notation and preliminary lemmas
$\phi^{*}(X)\geq\phi^{*}(\eta\mathbf{I})=\eta\phi^{*}(\mathbf{I})=\eta\mathbf{I}.$
@@
$\int_{\Omega}\left|\log\|(\varphi_{\omega}^{(\alpha)})^{*}\|_{\textnormal{op},1}\right|+\left|\log v((\varphi_{\omega}^{(\alpha)})^{*})\right|\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega)<\infty.$
@@
$\int_{\Omega}\left|\log\|\varphi_{\omega}^{(\alpha)}\|_{\textnormal{op},1}\right|+\left|\log v(\varphi_{\omega}^{(\alpha)})\right|\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega)<\infty.$
@@
We start with $\alpha=0$. Because $\varphi_{\omega}$ is CPTP, we have $\|\varphi_{\omega}\|_{\textnormal{op},1}=1$ and $v(\varphi_{\omega})=1$. For the adjoint, we will use the fact that $\varphi^{*}_{\omega}$ is then positive and unital. A lower bound is easily obtained: $\|\varphi^{*}_{\omega}\|_{\textnormal{op},1}\geq\|\varphi^{*}_{\omega}(\mathbf{I})\|_{1}/\|\mathbf{I}\|_{1}=\|\mathbf{I}\|_{1}/\|\mathbf{I}\|_{1}=1$. To obtain an upper bound, we note that $\|\varphi^{*}_{\omega}\|_{\textnormal{op},1}\leq d\|\varphi^{*}\|_{\textnormal{op},\infty}=d\|\varphi\|_{\textnormal{op},1}=d$. Finally, $\ln v(\varphi_{\omega}^{*})\in L^{1}$ by Assumption (A1).
@@
Hence, at fixed $\alpha\in{\mathbb{R}}$, Assumptions (A2) and (A3) guarantee that there exist random variables $Z^{(\alpha)}:\Omega\mapsto\mathbb{S}^{\circ}$ and $\tilde{Z}^{(\alpha)}:\Omega\mapsto\mathbb{S}^{\circ}$ satisfying the identities
@@
for $\mathbb{P}$-almost every $\omega$. This is the content of Theorem 1 in [MS22]. If, in addition, (A1) holds, then we have the following alternate expression fors the Lyapunov exponent $\lambda(\alpha)$ first seen in (2.5):
$\lambda(\alpha)=\int_{\Omega}\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega).$ (4.4)
@@
Noting that (4.3) implies
$\log\operatorname{tr}[\tilde{Z}_{\theta\omega}^{(\alpha)}\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]=\log\operatorname{tr}[\tilde{Z}_{\theta\omega}^{(\alpha)}Z_{\omega}^{(\alpha)}]+\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})],$
we arrive at a particularly useful variant of (4.4):
@@ 4.3 Simultaneous existence of limiting objects
$\frac{1}{2}\|Y-Y^{\prime}\|_{1}\leq\mathop{\mathrm{\mathstrut d}}\nolimits(Y,Y^{\prime})\leq\frac{1}{\min\operatorname{sp}(Y^{\prime})}\|Y-Y^{\prime}\|_{1}.$
@@
$\kappa(\alpha):=\inf_{N\in{\mathbb{N}}}\frac{1}{N}\int\log c(\Phi_{N}^{(\alpha)})\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}.$ (4.8)
By Lemmas 3.13–3.15 in [MS22], for each $\alpha\in{\mathbb{R}}$, the number $\kappa(\alpha)$ lies in the interval $[0,1)$ and there is a set of full $\mathbb{P}$-measure on which
@@
$\mathop{\mathrm{\mathstrut d}}\nolimits(\Phi_{-N;\omega}^{(\alpha)}\cdot Y,Z_{\omega}^{(\alpha)})\leq c(\Phi_{-N;\omega}^{(\alpha)})\quad\text{and}\quad\mathop{\mathrm{\mathstrut d}}\nolimits(\Phi_{N;\omega}^{(\alpha)}\cdot Y,\tilde{Z}_{\omega}^{(\alpha)})\leq c(\Phi_{N;\omega}^{(\alpha)}).$ (4.10)
@@
Therefore, the inequalities in (4.13) follow from the continuity bound of Lemma 4.7 below and an abstract convergence result for continuous submultiplicative families stated and proved in an appendix (Lemma A.3). ∎
@@
If (A3) holds, then there exists a set $\Omega_{0}$ of full $\mathbb{P}$-measure such that
$\left|c(\Phi_{n,\omega}^{(\alpha)})-c(\Phi_{n,\omega}^{(\beta)})\right|\leq 8|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega},$ (4.14)
@@
where due to the homogeneity of the expression, the projective actions can be replaced by linear actions. By (A3) and countable additivity of probability, there exists a full $\mathbb{P}$-measure set $\Omega_{0}$ with the property that, for all $\omega\in\Omega_{0}$, we have
@@
$\mathop{\mathrm{\mathstrut d}}\nolimits(\Phi_{n,\omega}^{(\beta)}\cdot X,\Phi_{n,\omega}^{(\alpha)}\cdot X)\leq\frac{1-\mathrm{e}^{-2|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}}}{1+\mathrm{e}^{-2|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}}}\leq 4|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}.$ (4.18)
@@
For fixed $\omega\in\Omega_{0}$ as in Lemmas 4.4 and 4.6 and $\alpha\in{\mathbb{R}}$, the proof of Lemma 3.12 in [MS22] can be followed line by line to deduce that $Z^{(\alpha)}_{\omega}$ and $\tilde{Z}^{(\alpha)}_{\omega}$ belong to $\mathbb{S}^{\circ}$ and satisfy (4.3) and (4.10) for all $Y\in\mathbb{S}$ and $N\in{\mathbb{N}}$. Indeed, the $\mathbb{P}$-almost sure properties that are used in that proof are those concerned by Lemmas 4.4 and 4.6. Finally, by $\theta$-invariance of $\mathbb{P}$ and countable additivity, taking

### v2

@@ 1 Introduction
$p^{(n)}(a_{1},\dotsc,a_{n})=\operatorname{tr}[(\psi_{n,a_{n}}\circ\dotsb\circ\psi_{1,a_{1}})(\rho)]$
@@ 2 Setup and main result
$M_{\omega,\rho}^{(n)}(\alpha)=\operatorname{tr}\left[\left(\varphi_{\theta^{n-1}\omega}^{(\alpha)}\circ\cdots\circ\varphi_{\omega}^{(\alpha)}\right)(\rho)\right],$ (2.3)
@@
For every fixed $\alpha$, the results of [MS22, PS23] will apply and tell us — among other things — that there is a $\mathbb{S}^{\circ}$-valued random variable $Z^{(\alpha)}$ that satisfies the cocycle relation
$\varphi_{\omega}^{(\alpha)}\left(Z_{\theta^{-1}\omega}^{\alpha}\right)=Z_{\omega}^{(\alpha)}\operatorname{tr}\left[\varphi_{\omega}^{(\alpha)}\left(Z_{\theta^{-1}\omega}^{\alpha}\right)\right]$
@@
For the derivation of the relation (2.7) from Theorem 1 in [PS23], see Lemma 4.3 below. These objects are at the heart of the proof of Theorem 2.3, which has of three main steps and is carried out in Section 4. We summarize the main steps of the proof here:
@@
In view of the specific form of the dependence of $\varphi_{\omega}^{(\alpha)}$ in $\alpha$, we use arguments from [MS22, PS23] to show that there is an event of full measure on which the limiting random variables $Z^{(\alpha)}$ (and analogous objects for adjoints) and the top Lyapunov exponent $\lambda(\alpha)$ exist simultaneously for every deformation parameter $\alpha\in{\mathbb{R}}$. This is the content of Lemmas 4.7 and 4.9, below.
@@ 3 Application to entropy production
Suppose that we have measurable maps $\omega\mapsto U_{\omega}\in\operatorname{U}({\mathbb{C}}^{d}\otimes{\mathbb{C}}^{m})$ and $\omega\mapsto\xi_{\omega}$, where $\xi_{\omega}$ is a Gibbs state on ${\mathbb{C}}^{m}$ with spectral decomposition
@@
using rank-one projectors, an increasing energy labeling $\epsilon\mapsto E_{\omega,\epsilon}$ and a positive inverse temperature $\beta_{\omega}$. The assumption that the eigenvalues of every corresponding Hamiltonian are all simple is made for ease of presentation. The argument could be easily adapted to Hamiltonians with degenerate eigenvalues, although if the degeneracy of eigenvalues depends non-trivially on $\omega$, some care would need to be taken to adjust the argument to account for the fact that the number of possible measurement outcomes is then random, e.g. by adding fictitious outcomes of probability 0.
Then, consider instruments labeled by pairs $(\epsilon,\epsilon^{\prime})\in\{1,2,\dotsc,m\}^{2}$ of indices for the spectral decomposition, that are of the form
@@
$f_{\omega}(\epsilon,\epsilon^{\prime})=\beta_{\omega}(E_{\omega,\epsilon^{\prime}}-E_{\omega,\epsilon})$
@@
$\Sigma_{\omega,n}((\epsilon_{1},\epsilon^{\prime}_{1}),(\epsilon_{2},\epsilon^{\prime}_{2}),\dotsc,(\epsilon_{n},\epsilon^{\prime}_{n}))=\sum_{j=0}^{n-1}\beta_{\theta^{j}\omega}(E_{{\theta^{j}\omega},\epsilon^{\prime}_{j}}-E_{{\theta^{j}\omega},\epsilon_{j}}),$
@@
and antiunitary involutions $\tau:{\mathbb{C}}^{d}\to{\mathbb{C}}^{d}$ and $\tau^{\prime}:{\mathbb{C}}^{m}\to{\mathbb{C}}^{m}$ such that
@@
    In the setup of Example 2.1.i, the map $\mathsf{T}:(\omega_{n})_{n\in{\mathbb{Z}}}\mapsto(\omega_{-n})_{n\in{\mathbb{Z}}}$ clearly satisfies the condition on the interplay of $\mathsf{T}$ and $\theta$, and the condition on the interplay of $\mathbb{P}$ and $\mathsf{T}$ is then naturally interpreted as a generalization of detailed balance for Markov chains. With $\omega\mapsto U_{\omega}\in\operatorname{U}({\mathbb{C}}^{d}\otimes{\mathbb{C}}^{m})$ and $\omega\mapsto\xi_{\omega}$ depending only on $\omega_{0}$, the conditions on $\tau$ and $\tau^{\prime}$ further simplify to $(\tau\otimes\tau^{\prime})U_{\omega}=U^{*}_{\omega}(\tau\otimes\tau^{\prime})$ and $\tau^{\prime}\xi_{\omega}=\xi_{\omega}\tau^{\prime}$. Writing $U_{\omega}=\exp(\mathrm{i}t_{\omega}H_{\omega})$, note that such $\tau$ and $\tau^{\prime}$ exist as complex conjugations if there exist orthonormal bases of ${\mathbb{C}}^{d}$ and ${\mathbb{C}}^{m}$ (used to construct an orthonormal basis of ${\mathbb{C}}^{d}\otimes{\mathbb{C}}^{m}$) with respect to which $\xi_{\omega}$ and $H_{\omega}$ have real matrix elements.
@@
Fix $\omega$ and suppose that $\xi_{\omega}=\tau^{\prime}\xi_{\mathsf{T}\omega}\tau^{\prime}$ as required in TRI. Comparing the spectral decompositions following the above labeling conventions, we find that $\pi_{\mathsf{\omega},\epsilon}=\tau^{\prime}\pi_{\mathsf{T}\omega,\epsilon}\tau^{\prime}$ for each $\epsilon$.
@@
$\sigma_{\omega,\rho,n}((\epsilon_{1},\epsilon^{\prime}_{1}),(\epsilon_{2},\epsilon^{\prime}_{2}),\dotsc,(\epsilon_{n},\epsilon^{\prime}_{n})):=\log\frac{p^{(n)}_{\omega,\rho}((\epsilon_{1},\epsilon^{\prime}_{1}),(\epsilon_{2},\epsilon^{\prime}_{2}),\dotsc,(\epsilon_{n},\epsilon^{\prime}_{n}))}{p^{(n)}_{\mathsf{T}(\theta^{n-1}\omega),\rho}((\epsilon^{\prime}_{n},\epsilon_{n}),(\epsilon^{\prime}_{n-1},\epsilon_{n-1}),\dotsc,(\epsilon^{\prime}_{1},\epsilon_{1}))}$ (3.1)
@@
$\Delta_{\rho}=\frac{1}{\inf\operatorname{sp}\rho},$
@@ 4.1 Notation and preliminary lemmas
$\phi^{*}(X)\geq\phi^{*}(\eta\mathbf{I})=\eta\phi^{*}(\mathbf{I})=\eta\mathbf{I},$
@@
$\int_{\Omega}\left|\log\|(\varphi_{\omega}^{(\alpha)})^{*}\|_{\textnormal{op},1}\right|+\left|\log v((\varphi_{\omega}^{(\alpha)})^{*})\right|\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega)<\infty$
@@
$\int_{\Omega}\left|\log\|\varphi_{\omega}^{(\alpha)}\|_{\textnormal{op},1}\right|+\left|\log v(\varphi_{\omega}^{(\alpha)})\right|\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega)<\infty$
@@
We start with $\alpha=0$. Because $\varphi_{\omega}$ is CPTP, we have $\|\varphi_{\omega}\|_{\textnormal{op},1}=1$ and $v(\varphi_{\omega})=1$. For the adjoint, we will use the fact that $\varphi^{*}_{\omega}$ is then positive and unital. A lower bound is easily obtained: $\|\varphi^{*}_{\omega}\|_{\textnormal{op},1}\geq\|\varphi^{*}_{\omega}(\mathbf{I})\|_{1}/\|\mathbf{I}\|_{1}=\|\mathbf{I}\|_{1}/\|\mathbf{I}\|_{1}=1$. To obtain an upper bound, we note that $\|\varphi^{*}_{\omega}\|_{\textnormal{op},1}\leq d\|\varphi^{*}\|_{\textnormal{op},\infty}=d\|\varphi\|_{\textnormal{op},1}=d$. Finally, $\log v(\varphi_{\omega}^{*})\in L^{1}$ by Assumption (A1).
@@
At fixed $\alpha\in{\mathbb{R}}$, Assumption (A2) and the bounds in (4.2) guarantee that there exist random variables $Z^{(\alpha)}:\Omega\mapsto\mathbb{S}^{\circ}$ and $\tilde{Z}^{(\alpha)}:\Omega\mapsto\mathbb{S}^{\circ}$ satisfying the identities
@@
for $\mathbb{P}$-almost every $\omega$. This is the content of Theorem 1 in [MS22]. If, in addition, (A1) and (A3) hold, then we have the following alternate expression for the Lyapunov exponent $\lambda(\alpha)$:
$\lambda(\alpha)=\int_{\Omega}\log\operatorname{tr}[(\varphi_{\omega}^{(\alpha)})^{*}(\tilde{Z}_{\theta\omega}^{(\alpha)})]\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega),$ (4.4)
@@
Lemma 4.3.
The following identities hold for the Lyapunov exponents:
$\lambda(\alpha)=\int_{\Omega}\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega),$ (4.6)
and
@@
Proof.
Note that the cocycle relations (4.3) imply that
$\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]=\log\operatorname{tr}[\tilde{Z}_{\theta\omega}^{(\alpha)}\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]-\log\operatorname{tr}[\tilde{Z}_{\theta\omega}^{(\alpha)}Z_{\omega}^{(\alpha)}]$ (4.8)
and
$\log\operatorname{tr}[(\varphi_{\omega}^{(\alpha)})^{*}(\tilde{Z}_{\theta\omega}^{(\alpha)})]$ $=\log\operatorname{tr}[(\varphi_{\omega}^{(\alpha)})^{*}(\tilde{Z}_{\theta\omega}^{(\alpha)})Z_{\theta^{-1}\omega}^{(\alpha)}]-\log\operatorname{tr}[\tilde{Z}_{\omega}^{(\alpha)}Z_{\theta^{-1}\omega}^{(\alpha)}]$
$=\log\operatorname{tr}[\tilde{Z}_{\theta\omega}^{(\alpha)}\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]-\log\operatorname{tr}[\tilde{Z}_{\omega}^{(\alpha)}Z_{\theta^{-1}\omega}^{(\alpha)}].$
Therefore,
$\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]-\log\operatorname{tr}[(\varphi_{\omega}^{(\alpha)})^{*}(\tilde{Z}_{\theta^{\omega}}^{(\alpha)})]=\log\operatorname{tr}[\tilde{Z}_{\omega}^{(\alpha)}Z_{\theta^{-1}\omega}^{(\alpha)}]-\log\operatorname{tr}[\tilde{Z}_{\theta\omega}^{(\alpha)}Z_{\omega}^{(\alpha)}].$ (4.9)
Noting that the right-hand side is of the form $g-g\circ\theta$ and that the terms $\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]$ and $\log\operatorname{tr}[(\varphi_{\omega}^{(\alpha)})^{*}(\tilde{Z}_{\theta^{\omega}}^{(\alpha)})]$ are in $L^{1}(\Omega,\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P})$ by Lemma 4.2, we conclude from Lemma A.5 below that
$\int_{\Omega}\log\operatorname{tr}[\varphi_{\omega}^{(\alpha)}(Z_{\theta^{-1}\omega}^{(\alpha)})]-\log\operatorname{tr}[(\varphi_{\omega}^{(\alpha)})^{*}(\tilde{Z}_{\theta^{\omega}}^{(\alpha)})]\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}(\omega)=0.$ (4.10)
The identity (4.6) follows from (4.4) and (4.10), and the identity (4.7) then follows from (4.6) and (4.8). ∎
@@ 4.3 Simultaneous existence of limiting objects
$\frac{1}{2}\|Y-Y^{\prime}\|_{1}\leq\mathop{\mathrm{\mathstrut d}}\nolimits(Y,Y^{\prime})\leq\frac{1}{\min\operatorname{sp}(Y^{\prime})}\|Y-Y^{\prime}\|_{1}$
@@
$\log\kappa(\alpha):=\inf_{N\in{\mathbb{N}}}\frac{1}{N}\int\log c(\Phi_{N}^{(\alpha)})\mathop{\mathrm{\mathstrut d}}\nolimits\!\mathbb{P}.$ (4.12)
By Lemmas 3.12–3.14 in [MS22], for each $\alpha\in{\mathbb{R}}$, the number $\kappa(\alpha)$ lies in the interval $[0,1)$ and there is a set of full $\mathbb{P}$-measure on which
@@
$\mathop{\mathrm{\mathstrut d}}\nolimits(\Phi_{-N;\omega}^{(\alpha)}\cdot Y,Z_{\omega}^{(\alpha)})\leq c(\Phi_{-N;\omega}^{(\alpha)})\quad\text{and}\quad\mathop{\mathrm{\mathstrut d}}\nolimits((\Phi_{N;\omega}^{(\alpha)})^{*}\cdot Y,\tilde{Z}_{\omega}^{(\alpha)})\leq c(\Phi_{N;\omega}^{(\alpha)}).$ (4.14)
@@
Therefore, the inequalities in (4.17) follow from the continuity bound of Lemma 4.8 below and an abstract convergence result for continuous submultiplicative families stated and proved in Appendix A (Lemma A.3). ∎
@@
There exists a set $\Omega_{0}$ of full $\mathbb{P}$-measure such that
$\left|c(\Phi_{n,\omega}^{(\alpha)})-c(\Phi_{n,\omega}^{(\beta)})\right|\leq 2|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}$ (4.18)
@@
where due to the homogeneity of the expression, the projective actions can be replaced by linear actions. By the definition of $F$ in (A3) and countable additivity of probability, there exists a full $\mathbb{P}$-measure set $\Omega_{0}$ with the property that, for all $\omega\in\Omega_{0}$, we have
@@
$\mathop{\mathrm{\mathstrut d}}\nolimits(\Phi_{n,\omega}^{(\beta)}\cdot X,\Phi_{n,\omega}^{(\alpha)}\cdot X)\leq\frac{1-\mathrm{e}^{-2|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}}}{1+\mathrm{e}^{-2|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}}}\leq|\beta-\alpha|\sum_{m=0}^{n-1}F_{\theta^{m}\omega}.$ (4.22)
@@
For fixed $\omega\in\Omega_{0}$ as in Lemmas 4.5 and 4.7 and $\alpha\in{\mathbb{R}}$, the proof of Lemma 3.12 in [MS22] can be followed line by line to deduce that $Z^{(\alpha)}_{\omega}$ and $\tilde{Z}^{(\alpha)}_{\omega}$ belong to $\mathbb{S}^{\circ}$ and satisfy (4.3) and (4.14) for all $Y\in\mathbb{S}$ and $N\in{\mathbb{N}}$. Indeed, the $\mathbb{P}$-almost sure properties that are used in that proof are those concerned by Lemmas 4.5 and 4.7. Finally, by $\theta$-invariance of $\mathbb{P}$ and countable additivity, taking